it takes richard semmler 3 hours to prepare a little league baseball field. it takes larry gilligan 7 hours to prepare the same field. how long will it take richard and larry working together to prepare a little league baseball field?
(simplify your answer. round to the one decimal place as needed)

Answer by richwmiller(17219)

1/7+1/3=1/x
3/21+7/21=1/x
10/21=1/x
21/10=x
2.1=x